OpenStudy (jaredstone4):

A battleship steams due east at 24 km/h. A submarine 4.0km away fires a torpedo that has a speed of 50 km/h. If the bearing of the ship as seen from the submarine is 20 degrees north of east... (a) In what direction should the torpedo be fired to hit the ship? (b) How long will it take the torpedo to reach the battleship?

OpenStudy (jaredstone4):

@ShailKumar

OpenStudy (jaredstone4):

On the test corrections, as hints I guess, it says "Set up position as a function of time functions for the battleship and the submarine, then set their final positions equal to each other." & "Use the distance that the torpedo must travel north in order to strike the ship, to come up with an expression for time."

OpenStudy (anonymous):



OpenStudy (anonymous):

Try to understand the problem. Let the time taken to hit the ship is t and the angle of projection of torpedo is \(\theta\) from east Horizontal motion: \(50 \cos \theta \times t = 24\times t + 4 \cos 20\) -------------(1) and vertical motion: \( 50 \sin \theta \times t = 4 \sin 20\) ---------------------(2) Solve these 2 equation for t and \(\theta\)

OpenStudy (anonymous):

Hope this helps...

OpenStudy (anonymous):